OpenStudy (anonymous):

Can someone show me how to figure this out? Find the coordinates of the midpoint of the segment whose endpoints are given. A (2, 3) and B (6, 11)

OpenStudy (anonymous):

take the average in each coordinate, aka add them up, divide by 2

OpenStudy (anonymous):

so 5 and 17 divide them by 2?

OpenStudy (anonymous):

\[\left(\frac{2+6}{2},\frac{3+11}{2}\right)\] is the first step

OpenStudy (anonymous):

Oh okay I see. Now what?

OpenStudy (anonymous):

compute what is \(\frac{2+6}{2}\) ?

OpenStudy (anonymous):

\[\frac{ 8 }{ 2 }\]

OpenStudy (anonymous):

ok, what is eight divided by two?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

sorry, forgot to simplify

OpenStudy (anonymous):

right, so first coordinate is 4

OpenStudy (anonymous):

how about the second coordinate?

OpenStudy (anonymous):

=7

OpenStudy (anonymous):

right so midpoint is \((4,7)\)

OpenStudy (anonymous):

Ahh. Thank you so much!